How many gallons of a plant food that is 6% nitrogen must be combined with another plant food that is 30% nitrogen to make 10 gal of a plant food that is 12% nitrogen?
Answer by Alan3354(69443)

Do it like this one:
------
I need 150 gallons of 10% solution but all i have is 5% and 20% solution. How much of the 5% and 20% solution do i need in order ti create 150 gallons of 10% solution?
===============
t = amount of 20%
f = amount of 5%
==========
t + f = 150
20t + 5f = 10*150